# ABScribe: Rapid Exploration of Multiple Writing Variations in Human-AI Co-Writing Tasks using Large Language Models

### v1

@@ 6. Discussion
In this work, we present the design and implementation of ABScribe which is composed of five key design elements: Variation Components, Hover Buttons, the Variation Accordion, AI Buttons, and AI Insert. Our comparative evaluation study shows that these elements reduce task workload and significantly enhance user perceptions of the revision process when managing multiple variations of text segments in human-AI co-writing tasks as compared to a familiar baseline editing system (Section 4).
We present six findings (F1-6, Sections 5.1 and 5.2) that provide evidence for the efficacy of our design elements. These findings offer insights into how writers utilize both chat-based and in-place Human-AI co-writing interfaces (Sections 2.3, 5.1). They also illustrate the affordances and limitations of these interface types for exploring multiple writing variations, their influence on the size, granularity, quantity, and diversity of variations, as well as the prompt style users take on.
6.1. Non-Linear Text Revision Control
In ABScribe, the Variation Component, and dynamically placed Hover Buttons provide an effective approach to managing text variations non-linearly. This offers an alternative to the linear text revision control features found in current editing interfaces and chat-based AI-assistants.
We find distinct influences on task workload (Section 5.2) and the style of prompts created by users between these two methods (F3-4, Section 5.1). Our interface is grounded in a non-linear nature of the writing and editing process (Sommers, 1980; Seow, 2002), offering a way for designers to support multiple fine-grained variations without overwhelming users (F2, Section 5.1). As our design for the Hover Buttons and Variation Components builds upon a familiar rich-text editing interface, they can potentially be integrated into existing document editors without major layout changes. This would enable writers to work more closely with LLM-based generative AI content, within the context of the surrounding text (F7, Section 5.2).
We also observed that the non linear approach affords advantages such as viewing, editing, and combining multiple variations together. Notably, participants W1 and W4 highlighted the usefulness of the non-linear text revision control even without AI (F6, Section 5.2), indicating the broader relevance of the Hover Button and Variation Component interface overlay for the wider range of text editors.
6.2. Scaffolding Prompts Focused Around Specific Writing Tasks
Our design for AI Buttons demonstrates how scaffolding LLM prompts around interface elements can influence user prompt-writing behavior. Specifically, the AI buttons, auto-generated after writers create a prompt for a text-segment variation, encourages writers to craft more direct, imperative prompts (F3, Section 5.1). It also made them reflect more on the revision process, shifting the focus from conversing with a chatbot to designing a button that represents their writing style for reuse across variations (F9, Section 5.2). We noticed a trade-off: while making prompts more reusable, it nudged users away from conversational prompts tied to a specific variation’s nuances (F4, Section 5.1). For example, they were more polite in the chat-based interface and direct when using the button scaffold. This highlights the importance of considering how different UI scaffolds influence the kinds of prompts users write.
These findings contribute to the body of research on encouraging writers’ experimentation of LLM output using modular LLM interfaces (Han et al., 2020b), and expands the scope of valuable design paradigms in text-editing, such as the use of reified text selections (Han et al., 2020a; Han, 2020; Beaudouin-Lafon and Mackay, 2000), to supporting the exploration of multiple variations in LLM-based revision-focused writing.
6.3. Moving Beyond Systematic Exploration to Systematic Evaluation of Variations
Our interface elements could be further developed to support the systematic evaluation of variations. While we want to avoid premature elimination of rough ideas (Greenberg and Buxton, 2008; Buxton, 2010), we also want to afford critical assessment of variations in relation to each other (Tohidi et al., 2006; Buxton, 2010). For instance, if we develop an extension that lets writers test different versions of the ABScribe draft containing a randomized subset of variations, they could link our interface to an open-source A/B testing framework like MOOClets (Reza et al., 2021), UpGrade (Murphy et al., [n. d.]) or Planout (Bakshy et al., 2014). This would allow writers to quantitatively evaluate which versions best meet their objectives based on specific metrics. A copywriter, for example, wanting to rapidly explore and assess different advertisement versions, can use ABScribe to explore draft variations, compare variations in the draft’s context using Hover Buttons, select a subset for further evaluation, and run randomized online experiments via an A/B testing framework. Moving beyond systematic exploration, which our current study covers, towards systematic evaluation is a logical next step. This opens a diverse design space for A/B authoring tools that simplify designing variations for evaluation, making it a promising avenue for future research.
Our design seeks to enhance the natural flow of writing and editing texts by supporting close collaboration between humans and AI, grounded in established theories on the revision process. As highlighted in Section 2.1, revising text is not merely about superficial changes (Fitzgerald, 1987), it delves into deeper subprocesses like idea formulation (Flower and Hayes, 1981a) and meaning discovery (Sommers, 1980). Skilled writers view revision as a cyclical and recursive process (Sommers, 1980; Seow, 2002), diving into the content repeatedly with different aims such as shaping a persuasive narrative (Sommers, 1980), playing with word choices and tone (Huff, 1983), and switching between various writing tasks during text revision (Faigley and Witte, 1981). Our interface provides a concrete example of affordances that editing interface designers can leverage to make the Human-AI collaboration in revisions more congruent with established revision theory.
In Section 2.1, we hypothesized that improving ease of use for handling multiple variations could help us effectively apply HCI design principles, affording the user greater freedom in experimenting with variations, thereby avoiding premature fixation on an initial idea (Tohidi et al., 2006; Jansson and Smith, 1991; Tohidi et al., 2006). As we found in F1 (5.1), the design of ABScribe helps writers to not commit to an initial idea too early, and instead, explore a greater quantity of variations before evaluating them, providing evidence for the applicability of parallel exploration in text revision.
6.4. Limitations
@@
Our work has two limitations to external validity, which are common in lab-based evaluation studies. First, our evaluation study was restricted to English writers. Although the underlying LLM, GPT-4, supports multiple languages (OpenAI, 2023), suggesting that our interface can extend beyond English, our study focused on English writing tasks. Without specifically studying how different languages influence task workload and user perceptions of the revision process, we cannot comment on the wider applicability of our tools, and would caution designers against directly applying our design without further evaluation. Second, our evaluation study was limited to a single 1.5-hour session with two guided writing tasks to ensure we could conduct our comparison in a controlled setting. We crafted realistic writing tasks that could be completed within the study’s timeframe, representing use cases from prior studies and commercial AI apps. Ideally, writers would select their own writing task and spend an extended period, possibly spanning several days or even a week, to revise and explore variations.
To address some of these concerns, we complemented the writing tasks with in-depth user interviews, allowing writers to reflect on and discuss the implications of our design beyond the study’s writing tasks. Several participants, such as W1, 3, 4, 11, and 12, expressed interest in using our tool for creative and academic writing, in settings beyond the use cases we explored.
@@ 7. Conclusion
In this work, we presented ABScribe, a human-AI co-writing interface built for swiftly exploring multiple writing variations using Large Language Models (LLMs).
Our interface is composed of an ensemble of five distinct elements: Variation Components, Variation Accordion, Hover Buttons, AI Buttons, and AI Insert. Collectively, these elements not only markedly decrease task workload but also bolster user perceptions of the revision process , in comparison to a popular AI-integrated editing workflow consisting of a rich text editor augmented with a chat-based AI assistant and the ability to insert AI generated content.
Our evaluation with writers (N=12) validate the efficacy of our design and offer insight into how writers leverage LLMs to explore variations, revealing a preference for non-linear over linear revision strategies, especially when engaging with a multitude of variations at finer granularity levels. We also found that scaffolding LLM use with task-focused UI components, like buttons, encouraged writers to create more generalizable prompts and use more direct, imperative language in prompt design. Our work informs HCI research on the design of Human-AI writing interfaces for the rapid exploration of writing variations.
@@ Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.References
@@
- bin (2013) 2013. Reinventing search with a new AI-powered Microsoft Bing and Edge, your copilot for the web. https://blogs.microsoft.com/blog/2023/02/07/reinventing-search-with-a-new-ai-powered-microsoft-bing-and-edge-your-copilot-for-the-web/.
- cha (2022) 2022. Introducing ChatGPT. https://openai.com/blog/chatgpt.
- PaL (2023) 2023. PaLM 2 Technical Report. https://ai.google/static/documents/palm2techreport.pdf.
- Bakshy et al. (2014) Eytan Bakshy, Dean Eckles, and Michael S Bernstein. 2014. Designing and deploying online field experiments. In Proceedings of the 23rd international conference on World wide web. 283–292.
- Baym et al. (2019) Nancy Baym, Limor Shifman, Christopher Persaud, and Kelly Wagman. 2019. INTELLIGENT FAILURES: CLIPPY MEMES AND THE LIMITS OF DIGITAL ASSISTANTS. AoIR Selected Papers of Internet Research 2019 (Oct. 2019). https://doi.org/10.5210/spir.v2019i0.10923
@@
- Bly (2007) Sara Bly. 2007. Fundamentals in HCI: Learning the Value of Consistency and User Models. (2007).
@@
- Buxton (2010) Bill Buxton. 2010. Sketching user experiences: getting the design right and the right design. Morgan kaufmann.
@@
- Cheung (2016) Yin Ling Cheung. 2016. Teaching writing. English language teaching today: Linking theory and practice (2016), 179–194.
@@
- Giray (2023) Louie Giray. 2023. Prompt Engineering with ChatGPT: A Guide for Academic Writers. Annals of Biomedical Engineering (2023), 1–5.
- Goldberg (2016) Natalie Goldberg. 2016. Writing down the bones: Freeing the writer within. Shambhala Publications.
@@
- Goodman et al. (2022) Steven M Goodman, Erin Buehler, Patrick Clary, Andy Coenen, Aaron Donsbach, Tiffanie N Horne, Michal Lahav, Robert MacDonald, Rain Breaw Michaels, Ajit Narayanan, et al. 2022. Lampost: Design and evaluation of an ai-assisted email writing prototype for adults with dyslexia. In Proceedings of the 24th International ACM SIGACCESS Conference on Computers and Accessibility. 1–18.
@@
- Han (2020) Han L Han. 2020. Designing Representations for Digital Documents. In Adjunct Proceedings of the 33rd Annual ACM Symposium on User Interface Software and Technology. 174–178.
- Han et al. (2020a) Han L Han, Miguel A Renom, Wendy E Mackay, and Michel Beaudouin-Lafon. 2020a. Textlets: Supporting constraints and consistency in text documents. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems. 1–13.
@@
- Harris (2017) Joseph Harris. 2017. Rewriting: How to do things with texts. University Press of Colorado.
- Hart (2006) Sandra G Hart. 2006. NASA-task load index (NASA-TLX); 20 years later. In Proceedings of the human factors and ergonomics society annual meeting, Vol. 50. Sage publications Sage CA: Los Angeles, CA, 904–908.
- Head et al. (2019) Andrew Head, Fred Hohman, Titus Barik, Steven M Drucker, and Robert DeLine. 2019. Managing messes in computational notebooks. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. 1–12.
- Hemingway (2014) Ernest Hemingway. 2014. Moveable feast: the restored edition. Simon and Schuster.
@@
- Jiang et al. (2022) Ellen Jiang, Kristen Olson, Edwin Toh, Alejandra Molina, Aaron Donsbach, Michael Terry, and Carrie J Cai. 2022. Promptmaker: Prompt-based prototyping with large language models. In CHI Conference on Human Factors in Computing Systems Extended Abstracts. 1–8.
- Kim et al. (2023) Tae Soo Kim, Yoonjoo Lee, Minsuk Chang, and Juho Kim. 2023. Cells, Generators, and Lenses: Design Framework for Object-Oriented Interaction with Large Language Models. (2023).
@@
- Lee et al. (2022a) Jaewook Lee, Jaylin Herskovitz, Yi-Hao Peng, and Anhong Guo. 2022a. ImageExplorer: Multi-Layered Touch Exploration to Encourage Skepticism Towards Imperfect AI-Generated Image Captions. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems. 1–15.
- Lee et al. (2022b) Mina Lee, Percy Liang, and Qian Yang. 2022b. Coauthor: Designing a human-ai collaborative writing dataset for exploring language model capabilities. In Proceedings of the 2022 CHI conference on human factors in computing systems. 1–19.
- Lee et al. (2022c) Mina Lee, Megha Srivastava, Amelia Hardy, John Thickstun, Esin Durmus, Ashwin Paranjape, Ines Gerard-Ursin, Xiang Lisa Li, Faisal Ladhak, Frieda Rong, et al. 2022c. Evaluating human-language model interaction. arXiv preprint arXiv:2212.09746 (2022).
@@
- MacArthur (2018) Charles A MacArthur. 2018. Evaluation and revision. Best practices in writing instruction 287 (2018).

### v3

@@ 6. Discussion
Writing plays a crucial role in many people’s lives, and the interfaces used for editing text can deeply influence the writing process. As LLMs increasingly blur the boundaries between human and AI writing capabilities (Zhou et al., 2022; Grace et al., 2018), the HCI community needs to explore interfaces that improve writers’ interactions with AI. In our work, we contributed an ensemble of five interface elements that significantly enhance user perceptions on AI-assisted revision (RQ1) by supporting rapid exploration and organization of multiple writing variations without overwhelming users (RQ2).
6.1. Design Implications
To interrogate the broader applicability of our results toward other contexts, and to characterize how ABScribe may relate to creativity support tools/authoring applications in general, we consider the following design implications:
D1: Emphasize Parallel Storage of Text Variations To Nudge Users Toward Working with a Larger Number of Increasingly Granular Ideas In line with HCI design principles that highlight the importance of investigating multiple alternatives (Buxton, 2010; Tohidi et al., 2006), we have developed ABScribe to facilitate the rapid exploration and organization of many text variations, which is a problem exacerbated by AI. Our findings suggest that parallel storage is a useful approach for inspiring writers to generate numerous ideas for granular sections of text, without becoming overly attached to any single variation. Although our implementation deals with text-segments, the challenge of organizing proliferating ideas through parallel storage extends beyond textual content, to other forms of creative content. As companies advance foundational models for manipulating various types of inputs — such as Google’s recently announced ‘natively multimodal’ Gemini model (Pichai and Hassabis, 2023), described as being adept at ‘reasoning seamlessly across text, images, video, audio, and code’ 666https://deepmind.google/technologies/gemini/ — designers can explore ways to extend our design to creativity support tools for other modalities. For example, an application for editing AI-generated images could implement an AI Modifier which accepts textual instructions and generates corresponding image variations, which are then organized within Variation Fields, contributing to HCI research on text-to-image AI editors (Brade et al., 2023; Liu, 2023; Liu and Chilton, 2022; Oppenlaender, 2022; Wang et al., 2023). The Popup Toolbar could be adapted to non-linear layer management in image editors (Chen et al., 2011; Barrett and Cheney, 2002), or applied in presentation software, allowing users to experiment with different headings and bullet points on a slide.
D2: De-Emphasize Parallel Storage to Support Editing Forms that Are Better Suited to Linear Representations of Content. Our findings also suggest that while parallel storage nudges users towards exploring multiple text variations, it remains important to cater to linear editing preferences. Many participants preferred to view previous variations side-by-side or in-sequence when producing a new variation. For example, W6 noted their habit of linearly organizing text variations for side-by-side comparison, which seemed less intuitive in ABScribe. Although our design does not prevent the sequential pasting of text blocks, it subtly discourages linear editing. This insight underscores the potential benefits of a Variation Sidebar as a supplementary tool for linear viewing of variations, complementing the Popup Toolbar’s parallel functionality. A more complex Variation Sidebar that allows for reordering and editing inside the accordion could help with linear revisions. The Variation Fields could also be changed to allow for both parallel and linear editing, such as through incorporating an ’explode/contract’ button in the Popup Toolbar that allow users to switch between a linear, exploded display of variations for easier comparison and a more compact, collapsed view within the Variation Field. Together, D1 and D2 contribute to existing discussions surrounding linear versus non-linear representations of versions history across various domains (Head et al., 2019; O’Sullivan, 2009; Chen et al., 2011; Plaice and Wadge, 1993; Zhang et al., 2023; Tichy, 1982).
D3: Scaffold Prompts Within Discrete Interface Elements to Encourage Action-Oriented Prompt-Writing. If the goal for designers is to guide users towards crafting more succinct and direct prompts, incorporating scaffolding around specific interface elements signifying actions can be effective. This approach is evident in the buttons created by AI Modifiers. In contrast, in free-form conversational user interfaces, we saw a tendency for users to adopt a more polite and verbose style. For example, participant W8 would say, “Can you please make this variation shorter?” accompanied by a pasted text snippet, rather than simply selecting a Variation Field and requesting, “Make it short.” Of course, there are contexts where encouraging a polite and detailed interaction with AI agent is beneficial, e.g., in training scenarios where a customer service representative practices dialogues with an AI. However, in situations focused on productivity tasks, such as composing an email, steering users towards a direct prompt-writing styles can help. Our findings illustrate how interface elements can noticeably influence prompt-writing style, akin to previous work on, such as Dang et al.’s investigation of diegetic and non-diegetic prompts (Dang et al., 2023). It also adds to prior work using objects to represent AI prompts (Kim et al., 2023a) and text segments (Han et al., 2020a; Beaudouin-Lafon and Mackay, 2000).
D4: Reduce User Effort by Turning AI Prompts into Customizable Toolkit for Digital Tasks. AI Modifiers in our interface highlights the potential of LLM prompts as tailored sets of reusable user actions, resembling a bespoke Swiss Army knife. For instance, users who frequently need to condense, translate, and proofread text can effortlessly create AI modifiers for these tasks, facilitating reuse and minimizing effort. Previous studies have shown that users may hesitate to write prompts when given multiple AI suggestions (Dang et al., 2023), underscoring the importance of reducing prompt-writing effort. Our design addresses aspects of this problem.
Unlike traditional menu buttons which have predictable outcomes, AI Modifiers can yield unexpected results due to the inherent unpredictability of AI models (Yampolskiy, 2020, 2019; Michel, 2020; Hauhio et al., 2023). Prior research on AI-assisted editing suggests that users may find imperfect AI outputs useful (Yuan et al., 2022), given that they have the ability to edit. Thus, while using AI prompts as personalized toolkits is beneficial, designers will have to ensure that users can easily refine both the AI’s imperfect output and the unpredictable instructions encapsulated by the AI modifier.
The current implementation of AI Modifiers as labeled buttons is only one of many possible approaches. Alternative user experiences could include AI-generated icons, options within a nested menu, or even simplified voice commands, for example, a user might initially issue a detailed voice command and then receive a suggestion for a condensed version for future use. Our design could be also be extended to give users more control over AI features, such as adjusting the ’temperature’ to modulate predictability.
6.2. Limitations and Future Work
A noteworthy limitation of our work is that we constructed the baseline interface rather than opting to use an existing system in our evaluation. We chose to construct the baseline because early pilots of our evaluation study revealed that utilizing a real-world system as a baseline (in our case, ChatGPT and GoogleDocs) introduced several logistical challenges that made it difficult to control and run the study. For instance, to ensure ChatGPT used the same underlying LLM model, we had to provide participants with login details to a paid ChatGPT account, potentially exposing credit card details. When using Google Docs, participants would resort to using editing features outside the focus of our research, such as user comments. Our choice of constructing a baseline is in alignment with prior HCI research on AI writers (Yuan et al., 2022). However, this may have affected the ecological validity of our results. Future studies could compare ABScribe with real-world AI-assisted writing workflows.
@@ 6.2. Limitations and Future Work
There are also limitations to external validity common in lab-based evaluation studies. First, our evaluation study was restricted to English writers. Although the underlying LLM, GPT-4, supports multiple languages (OpenAI, 2023), our study focused on English writing tasks. Therefore, we would caution designers against directly applying our design without further evaluation. Second, our evaluation study was limited to a single 1.5-hour session with two guided writing tasks to ensure we could conduct our comparison in a controlled setting. Future work could investigate how writers select their own writing task and revise text over extended periods, possibly spanning several days.
A potential risk with any generative AI tool is that the AI may produce hallucinated or confabulated content (Salvagno et al., 2023; Emsley, 2023; Smith et al., 2023; Ji et al., 2023). ABScribe’s AI Drafter and AI Modifiers both present a risk of introducing hallucinations into the text, which becomes especially important to be cognizant of when writing in high-stakes domains or in educational settings. We carefully named our features to account for this. With the AI Drafter, the tool is specifically referred to as a drafter, not a writing companion, to nudge the user toward thorough revision of what it produces. The AI Modifier rephrases existing text written by the author, which provides the generative AI with a baseline grounded in the author’s text, which is expected to be fact-checked.
Finally, it would be useful to implement logging features into ABScribe to concrete measures from the tool, such as: How many words did the subjects actually write; how many words were written by the LLM; what percentage of the resulting text was from the LLM vs. the human; how many times did the subjects interact with the various features in the ABScribe UI? We did not collect these metrics as our focus was on addressing task workload and user perceptions. Future studies could investigate the relative contributions of human and AI in co-written documents.
@@ 7. Conclusion
In this work, we presented ABScribe, a human-AI co-writing interface built for swiftly exploring multiple writing variations using Large Language Models (LLMs). Our design only markedly decreasees task workload but also bolsters user perceptions of the revision process , in comparison to a baseline chat-based AI editing workflow. We evaluated our design with writers (N=12) to validate the efficacy of ABScribe and offer insight into how writers leverage LLMs to explore variations, revealing a preference for non-linear over linear revision strategies, especially when engaging with a multitude of variations at finer granularity levels. We also found that scaffolding LLM use with task-focused UI components, like buttons, encouraged writers to create more generalizable prompts and use more direct, imperative language in prompt design. Our work informs HCI research on the design of Human-AI writing interfaces for the rapid exploration of writing variations.
@@ Acknowledgements.
We are grateful to the writers who participated in this work. We would also like to thank Siqi (Jessica) Zhang and Divya Gupta for their assistance with the transcription of interviews, and Steven Tohme for his help in open-sourcing the codebase. This work was partially supported by grants from the Office of Naval Research (N00014-21-1-2576), the Natural Sciences and Engineering Research Council of Canada (RGPIN-2019-06968), and the National Science Foundation (2209819).References
@@
- Barrett and Cheney (2002) William A Barrett and Alan S Cheney. 2002. Object-based image editing. In Proceedings of the 29th annual conference on Computer graphics and interactive techniques. Association for Computing Machinery, New York, NY, USA, 777–784.
- Baym et al. (2019) Nancy Baym, Limor Shifman, Christopher Persaud, and Kelly Wagman. 2019. INTELLIGENT FAILURES: CLIPPY MEMES AND THE LIMITS OF DIGITAL ASSISTANTS. AoIR Selected Papers of Internet Research 2019 (Oct. 2019), 1–5. https://doi.org/10.5210/spir.v2019i0.10923
@@
- Biermann et al. (2022) Oloff C. Biermann, Ning F. Ma, and Dongwook Yoon. 2022. From Tool to Companion: Storywriters Want AI Writers to Respect Their Personal Values and Writing Strategies. In Proceedings of the 2022 ACM Designing Interactive Systems Conference (Virtual Event, Australia) (DIS ’22). Association for Computing Machinery, New York, NY, USA, 1209–1227. https://doi.org/10.1145/3532106.3533506
- Bly (2007) Sara Bly. 2007. Fundamentals in HCI: Learning the Value of Consistency and User Models. In HCI Remixed: Reflections on Works That Have Influenced the HCI Community. The MIT Press, Cambridge, MA, USA. https://doi.org/10.7551/mitpress/7455.003.0012 arXiv:https://direct.mit.edu/book/chapter-pdf/2290056/9780262256070_cah.pdf
@@
- Buschek et al. (2021) Daniel Buschek, Martin Zürn, and Malin Eiband. 2021. The impact of multiple parallel phrase suggestions on email input and composition behaviour of native and non-native english writers. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New York, NY, USA, 1–13.
- Buxton (2010) Bill Buxton. 2010. Sketching user experiences: getting the design right and the right design. Morgan kaufmann, Boston.
- Chen et al. (2011) Hsiang-Ting Chen, Li-Yi Wei, and Chun-Fa Chang. 2011. Nonlinear revision control for images. ACM Transactions on Graphics (TOG) 30, 4 (2011), 1–10.
@@
- Cheung (2016) Yin Ling Cheung. 2016. Teaching Writing. Springer International Publishing, Cham, 179–194. https://doi.org/10.1007/978-3-319-38834-2_13
- Chung et al. (2022) John Joon Young Chung, Wooseok Kim, Kang Min Yoo, Hwaran Lee, Eytan Adar, and Minsuk Chang. 2022. TaleBrush: Sketching stories with generative pretrained language models. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New York, NY, USA, 1–19.
@@
- Dang et al. (2023) Hai Dang, Sven Goller, Florian Lehmann, and Daniel Buschek. 2023. Choice over control: How users write with large language models using diegetic and non-diegetic prompting. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New York, NY, USA, 1–17.
@@
- Devlin et al. (2018) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2018. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. arXiv:1810.04805 http://arxiv.org/abs/1810.04805
@@
- Emsley (2023) Robin Emsley. 2023. ChatGPT: these are not hallucinations–they’re fabrications and falsifications. Schizophrenia 9, 1 (2023), 52.
@@
- Gero et al. (2022) Katy Ilonka Gero, Vivian Liu, and Lydia Chilton. 2022. Sparks: Inspiration for Science Writing using Language Models. In Proceedings of the 2022 ACM Designing Interactive Systems Conference (ustralia) (DIS ’22). Association for Computing Machinery, New York, NY, USA, 1002–1019. https://doi.org/10.1145/3532106.3533533
- Gero et al. (2023) Katy Ilonka Gero, Tao Long, and Lydia B Chilton. 2023. Social dynamics of AI support in creative writing. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New York, NY, USA, 1–15.
- Giray (2023) Louie Giray. 2023. Prompt Engineering with ChatGPT: A Guide for Academic Writers. Annals of Biomedical Engineering 51 (06 2023), 3. https://doi.org/10.1007/s10439-023-03272-4
- Goldberg (2016) Natalie Goldberg. 2016. Writing down the bones: Freeing the writer within. Shambhala Publications, 2129 13th St, Boulder, Colorado.
@@
- Goodman et al. (2022) Steven M Goodman, Erin Buehler, Patrick Clary, Andy Coenen, Aaron Donsbach, Tiffanie N Horne, Michal Lahav, Robert MacDonald, Rain Breaw Michaels, Ajit Narayanan, et al. 2022. Lampost: Design and evaluation of an ai-assisted email writing prototype for adults with dyslexia. In Proceedings of the 24th International ACM SIGACCESS Conference on Computers and Accessibility. Association for Computing Machinery, New York, NY, USA, 1–18.
- Google (2023) Google. 2023. PaLM 2 Technical Report. https://ai.google/static/documents/palm2techreport.pdf.
- Grace et al. (2018) Katja Grace, John Salvatier, Allan Dafoe, Baobao Zhang, and Owain Evans. 2018. When will AI exceed human performance? Evidence from AI experts. Journal of Artificial Intelligence Research 62 (2018), 729–754.
@@
- Han (2020) Han L Han. 2020. Designing Representations for Digital Documents. In Adjunct Proceedings of the 33rd Annual ACM Symposium on User Interface Software and Technology. Association for Computing Machinery, New York, NY, USA, 174–178.
- Han et al. (2020a) Han L Han, Miguel A Renom, Wendy E Mackay, and Michel Beaudouin-Lafon. 2020a. Textlets: Supporting constraints and consistency in text documents. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems. Association for Computing Machinery, New York, NY, USA, 1–13.
@@
- Harris (2017) Joseph Harris. 2017. Rewriting: How to do things with texts. University Press of Colorado, Boulder, CO, USA.
- Hart (2006) Sandra G. Hart. 2006. Nasa-Task Load Index (NASA-TLX); 20 Years Later. Proceedings of the Human Factors and Ergonomics Society Annual Meeting 50, 9 (2006), 904–908. https://doi.org/10.1177/154193120605000909 arXiv:https://doi.org/10.1177/154193120605000909
- Hauhio et al. (2023) Iikka Hauhio, Anna Kantosalo, Simo Linkola, and Hannu Toivonen. 2023. The Spectrum of Unpredictability and its Relation to Creative Autonomy. In Proceedings of the 14th International Conference on Computational Creativity. Association for Computational Creativity (ACC), Spain, 148–152. https://computationalcreativity.net/iccc23/ International Conference on Computational Creativity, ICCC ; Conference date: 19-06-2023 Through 23-06-2023.
- Head et al. (2019) Andrew Head, Fred Hohman, Titus Barik, Steven M. Drucker, and Robert DeLine. 2019. Managing Messes in Computational Notebooks. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3290605.3300500
- Hemingway (2014) Ernest Hemingway. 2014. Moveable feast: the restored edition. Simon and Schuster, New York, NY, USA.
@@
- Ji et al. (2023) Ziwei Ji, Nayeon Lee, Rita Frieske, Tiezheng Yu, Dan Su, Yan Xu, Etsuko Ishii, Ye Jin Bang, Andrea Madotto, and Pascale Fung. 2023. Survey of hallucination in natural language generation. Comput. Surveys 55, 12 (2023), 1–38.
- Jiang et al. (2022) Ellen Jiang, Kristen Olson, Edwin Toh, Alejandra Molina, Aaron Donsbach, Michael Terry, and Carrie J Cai. 2022. PromptMaker: Prompt-based Prototyping with Large Language Models. In Extended Abstracts of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI EA ’22). Association for Computing Machinery, New York, NY, USA, Article 35, 8 pages. https://doi.org/10.1145/3491101.3503564
- Kim et al. (2023a) Tae Soo Kim, Yoonjoo Lee, Minsuk Chang, and Juho Kim. 2023a. Cells, Generators, and Lenses: Design Framework for Object-Oriented Interaction with Large Language Models. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (San Francisco, CA, USA) (UIST ’23). Association for Computing Machinery, New York, NY, USA, Article 4, 18 pages. https://doi.org/10.1145/3586183.3606833
- Kim et al. (2023b) Yewon Kim, Mina Lee, Donghwi Kim, and Sung-Ju Lee. 2023b. Towards Explainable AI Writing Assistants for Non-native English Speakers.
@@
- Lee et al. (2022a) Mina Lee, Percy Liang, and Qian Yang. 2022a. Coauthor: Designing a human-ai collaborative writing dataset for exploring language model capabilities. In Proceedings of the 2022 CHI conference on human factors in computing systems. Association for Computing Machinery, New York, NY, USA, 1–19.
- Lee et al. (2022b) Mina Lee, Megha Srivastava, Amelia Hardy, John Thickstun, Esin Durmus, Ashwin Paranjape, Ines Gerard-Ursin, Xiang Lisa Li, Faisal Ladhak, Frieda Rong, et al. 2022b. Evaluating human-language model interaction.
@@
- Lewis and Wharton (1997) Clayton Lewis and Cathleen Wharton. 1997. Chapter 30 - Cognitive Walkthroughs. In Handbook of Human-Computer Interaction (Second Edition) (second edition ed.), Marting G. Helander, Thomas K. Landauer, and Prasad V. Prabhu (Eds.). North-Holland, Amsterdam, 717–732. https://doi.org/10.1016/B978-044481862-1.50096-0
@@
- Liu and Chilton (2022) Vivian Liu and Lydia B Chilton. 2022. Design Guidelines for Prompt Engineering Text-to-Image Generative Models. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 384, 23 pages. https://doi.org/10.1145/3491102.3501825
- Long et al. (2023) Tao Long, Dorothy Zhang, Grace Li, Batool Taraif, Samia Menon, Kynnedy Simone Smith, Sitong Wang, Katy Ilonka Gero, and Lydia B Chilton. 2023. Tweetorial Hooks: Generative AI Tools to Motivate Science on Social Media.
- MacArthur (2018) Charles A MacArthur. 2018. Evaluation and revision. Best practices in writing instruction 287 (2018), 23 pages.
@@
- Mahatody et al. (2010) Thomas Mahatody, Mouldi Sagar, and Christophe Kolski. 2010. State of the art on the cognitive walkthrough method, its variants and evolutions. Intl. Journal of Human–Computer Interaction 26, 8 (2010), 741–785.